Science and medicine. Find the time required for an object to fall to the ground from a building that is 1400 ft high. (Use the formula t=1/4Vs_) I'm using the V as a square root symbol.

Thank you.

Answer by Fermat(127)

The formula for the time taken to fall freely under gravity through a distance s is given by,
s = (1/2)gtē
Since you are workng in imperial units (feeet, pounds seconds), then g = 32 ft/sē
So,
s = (1/2)32tē
s = 16tē
tē = s/16
t = √(s)/4
==========
Since s = 1400,
t = √(1400)/4
t = 37.417/4
t = 9.354 s
===========